OpenStudy (holliday4u):

What is the equation of a circle with center (1, -4) and radius 2?

TheSmartOne (thesmartone):

\(\sf\Large (x-h)^2 + (y-k)^2 = r^2\) Where (h,k) is the center of the circle, and r is the radius.

OpenStudy (holliday4u):

thanks

TheSmartOne (thesmartone):

Also, welcome to OpenStudy! :)

OpenStudy (holliday4u):

(x-1)^2+(y--4)^2=2^2

TheSmartOne (thesmartone):

correct, and y--4 can be simplified to y + 4

OpenStudy (holliday4u):

yay i got it right thanks!